OpenStudy (anonymous):

Stacey has a jar that contains 21 green skittles and 30 yellow skittles. She withdraws one skittle and notes the color. She does not return it to the jar. Stacey then withdraws another skittle. Find the probability that the 1st skittle is green and the next skittle is yellow.

OpenStudy (anonymous):

and in this case means multiply. probability that first skittle is green is 21/51, probability that second is yellow GIVEN that first one is green is 30/50 so you get \[\frac{21}{51}\times \frac{30}{50}\]

OpenStudy (anonymous):

or \[\frac{63}{155}\] if you prefer